# Dressed to Gamble: How Poker Drives the Dynamics of Wearables and Visits on Decentraland’s Social Virtual World

### v2

@@ 3.1 Fashion in Virtual Worlds
Fashion in social virtual worlds has been a subject of research for years, especially in Second Life, which has been active since 2003 and introduced the concept of customizable virtual spaces and avatars Bardzell et al. (2010). Despite criticisms that it did not live up to public expectations Rodriguez Sanchez & Garcia-Badell (2022), Second Life maintains a robust economy centered on avatar customization, with many users becoming in-world fashion designers, creating clothing and accessories for avatars Boellstorff (2015).
@@
Recently, the fashion industry’s attention has shifted towards popular MMO games, particularly through the creation of branded skins—cosmetic options for characters, with an increasing interest in newer blockchain-based virtual worlds Rodriguez Sanchez & Garcia-Badell (2022). Notably, several prestigious luxury brands are venturing into NFT collectibles Joy et al. (2022), participating in events such as the Metaverse Fashion Week or designing wearables for Decentraland Gonzalez (2022), as well as retailing and engaging in other kinds of customer engagement within Decentraland’s social virtual world Goldberg et al. (2023).
In fact, Decentraland’s wearables are mainly distributed for free to boost engagement in other cryptoasset or Metaverse initiatives, with only a minor portion of them (3.4%) being sold on the platform’s marketplace Trujillo & Bacciu (2023). Besides wearables, other kinds of NFTs available within the virtual world (e.g., tickets, memberships) are likewise transferred for free as part of promotional activities instead of being part of a sale transaction Goldberg et al. (2023). Perhaps unsurprisingly, when sold in Decentraland’s marketplace, the price of wearables is largely influenced by their pre-established rarity Trujillo & Bacciu (2023), which escalates even further for unique items —i.e., having a set minting limit of one. Incidentally, over half of all wearable creators in Decentraland typically author only a single collection (mostly with a single item), but as mentioned in the Introduction, a specific account related to Decentral Games stands out having published circa 80 collections with more than 600 items total in a year and a half of activity Trujillo & Bacciu (2023). These wearables are the only way to get access to the games played on the company’s casinos within Decentraland, with the primary offering being ICE Poker, a virtual poker gaming experience.
3.2 Poker and Online Gambling
@@
The chance-based nature of the game of poker appeals to a wide audience of casual players, while the element of skill, particularly evident over many hands, entices professional players Fiedler & Rock (2009). The introduction of Internet to gambling activities has provided easy access for the general public, round-the-clock access, shorter intervals between bets, and instant reinforcements Griffiths & Barnes (2008). After 2003, online poker experienced tremendous growth, with a total market value reaching 3 to 4 billion USD by 2010, with approximately 5.5 million players participating in virtual poker rooms across the globe Fiedler & Wilcke (2011). Among these players, a subset has emerged as professionals, especially adept at multitabling —simultaneously playing multiple games to further reduce the element of chance. As reliance on poker earnings grew, these players developed a strong extrinsic motivation to play, even if not entirely removed from the gameplay. Further, it has been found that playing for real money strongly influences both game dynamics and player experiences in poker games Zaman et al. (2014).
@@
Notably, professional gamblers share psychological traits with problem gamblers, highlighting the complex interplay between skill and risk Newall & Talberg (2023). Problem gamblers, often drawn to low-stakes gambling, may develop distorted risk perceptions and ultimately underestimate financial risks in general Armstrong et al. (2018). Interestingly, players are socially motivated to gamble by the virtual communities they form, while these same communities also serve as safeguards against excessive behavior Sirola et al. (2021); Savolainen et al. (2022). Online poker’s unique subculture —which includes websites, forums, and chat rooms from all around the world— fosters cooperative behavior among players through the sharing of information and strategies. Nonetheless, the game’s inherent competitiveness persists and is reflected in the hierarchical structures within these global communities O’Leary & Carroll (2013).
Despite online gambling being a global phenomenon, its legality varies across jurisdictions, thereby posing significant challenges for both players and operators as they attempt to determine whether a specific game is legal to play or not Zborowska et al. (2012). The advent of blockchain technology has ignited enthusiasm due to its potential to improve fairness and transparency in the sector, promising decentralized and autonomous solutions for regulators and operators alike. However, at the same time it ultimately adds complexity to the legal framework of online gambling, due in great part to the current regulatory uncertainties of cryptocurrencies and other blockchain-based assets Gainsbury & Blaszczynski (2017).
3.3 Crypto Asset Trading
@@
Cryptocurrency trading has emerged as a disruptive force in the financial landscape since the inception of Bitcoin in 2008. While market capitalizations have reached trillions of dollars, the trading environment remains characterized by high volatility, often leading to substantial crashes followed by significant rebounds Fang et al. (2022); OECD (2022). NFTs entered the scene in 2014, gaining widespread attention in early 2021 due to very high-profile sales. Digital art, music albums, and even tweets and sports highlights have been registered as unique crypto assets and subsequently sold for millions of dollars each, causing considerable hype Houser & Holden (2022). NFT users exhibit a profound enthusiasm for the foundational technologies that facilitate unique avenues for creative expression and the development of innovative content creation business models. However, these positive sentiments are counterbalanced by the contentious matter of digital content ownership, the prevalence of low-quality NFTs, the risk of scams, potential money laundering activities, and the evolving framework of legal measures Sharma et al. (2022).
@@
The literature also shows an overlap of personality and demographic traits among problem gamblers and cryptocurrency traders Johnson et al. (2023); Delfabbro et al. (2021b). In both cryptocurrency trading and gambling, decisions are often based on limited information, short-term motives for gain, and with high uncertainty; hence, people attracted to latter are likely to be attracted to the former Delfabbro et al. (2021a). Numerous operators in the crypto gambling sector provide access to a variety of betting options, games, and online casinos. However, these platforms often pose a significant risk as they are easily accessible to minors and vulnerable individuals, and typically offer minimal to no consumer protection features. The legal landscape for these platforms is also a concern, as they remain largely unregulated in many countries Andrade et al. (2023).
Gambling aside, the legality of trading both cryptocurrencies and NFTs also varies greatly across jurisdictions. Some countries embrace these digital assets, while others impose strict regulations. Factors include tax treatment, anti-money laundering compliance, and investor protection. Researchers and policymakers continue to grapple with these complexities, seeking a balance between industry innovation and safeguarding the public Yeoh (2017).
3.4 Play-to-Earn Games
The emergence of blockchain technologies has also given rise to a new form of monetized online gaming known as “Play-to-Earn” (P2E) games. In P2E games, players receive monetary rewards for their gameplay, mainly in the form of the game’s specific cryptocurrency. Axie Infinity —a Pokémon-style battle game with collectible axolotl-like creatures— is arguably one of the most well-known examples, making use of cryptocurrencies and NFTs for its in-game economy. Its popularity peaked in 2021, particularly in the Philippines and later Venezuela, where it represented a viable source of earnings for low-income people during the height of the pandemic De Jesus et al. (2022).
@@ 3.4 Play-to-Earn Games
However, extrinsically motivated gameplay, i.e., mainly driven by external rewards such as financial gain, can reduce intrinsic motivation and consequently have negative implications for player experience Delic & Delfabbro (2024); Hong et al. (2023). A survey conducted with a group of Axie Infinity players in 2020–21 has shown that, while some players enjoyed social interactions, overall intrinsic motivation tended to be low, causing high amounts of stress and unwanted tension to reach daily quotas De Jesus et al. (2022). This assessment was later corroborated by a study that analyzed online conversations about the game Delic & Delfabbro (2024).
@@
Interestingly, part of the initial success of Axie Infinity was due to delegation models that emerged organically within the community, called scholarships De Jesus et al. (2022). These models allow players to share their knowledge or delegate their gameplay (and assets) to others, in exchange for a share of the rewards. Nonetheless, scholarship models and related delegation mechanisms within P2E games are potentially exploitative. While P2E games offer players from less-advantaged countries the prospect of earning income above the national average wage, players might find themselves participating in a scheme that provides low-cost labor for wealthier actors who hold valuable NFTs Delic & Delfabbro (2024). Curiously, ICE Poker integrates by design some elements of P2E games, offering the possibility to earn tokens just by participating in certain events, and a delegation mechanism similar to the community-based scholarships in Axie Infinity.
4 Methods
For our study, we focus on the two years after the launch of Decentraland wearables on Polygon, that is, from June 2021 to May 2023. We excluded the short-lived first version of wearables because these were already deprecated by the launch of ICE Poker. In the following, we describe the overall approach to answer our RQs and then we detail how we performed the respective data collection.
4.1 Overall approach
@@
For RQ1 (wearable transfers) we first retrieved all wearables created during the period of study, marking those created by DG based on metadata from Decentraland. We then retrieved from the Polygon blockchain all the transfers related to all wearables, as well as their sales made in the most prominent marketplaces. We then analyzed the proportion between DG wearables and the rest in terms of the number of transfers as well as sales volume (both in number and USD). We also analyzed the share of transfers related to upgrades, which we retrieved from the respective log events on Polygon.
@@
For RQ2 (in-world visits to DG poker casinos), we first collected in-world tracking data of users, available only from 12 July 2022. Hence, we analyzed the last 10 months of the two-year period of interest, in terms of the number of accounts and time spent visiting the parcels in which DG poker casinos are located, compared to the rest of Decentraland’s map.
@@
Our descriptive and statistical analyses mostly concern proportions, i.e., the share of wearables or land associated with DG compared with the rest of Decentraland. For RQ1, this is either for the whole two-year period or on a monthly basis therein. For RQ2, this is either for a ten-month period or on a daily basis, as we have a narrower time frame but a finer time resolution for data of in-world visits. To test for significant differences in proportions we used two-sided tests with significance level $\alpha=0.05$. When analyzing samples we used $\chi^{2}$ tests, either goodness-of-fit test or association test according to the case; when analyzing populations we used a binomial test; for both kinds of test we use relative risk (RR) to measure the effect size. For daily in-world visits, we also used iterative STL (i.e., seasonal and trend decomposition using LOESS) to analyze the time series, including trend and seasonal strengths (from 0 to 1).
4.2 Data collection
@@
For wearable metadata (e.g., collection, category, creator), we used several of the openly available Decentraland’s query endpoints on TheGraph, a service for querying blockchain networks, and whose definitions are open source. [^1] https://github.com/decentraland For transfers (including mints), we collected the respective transactions from Polygon via the blockchain web service Alchemy. For wearable sales, we collected data from two different sources: 1) Decentraland’s Marketplace for internal sales, using the aforementioned TheGraph query endpoints; and 2) OpenSea, the largest overall NFT marketplace and main means for external sales, as well as the initial default marketplace for DG wearables.
@@
We also retrieved from Yahoo! Finance the price of the few different cryptocurrencies used for the collected wearable sales during the two-year period of study. To normalize the price of sales, we first converted it to USD, then we adjusted it for inflation on a monthly basis as of December 2023, based on the Consumer Price Index (CPI) by the Bureau of Labor Statistics of the United States. Price changes of these cryptocurrencies also serve to gauge the interest on the studied platforms. Indeed, as can be seen in Figure 4, the ICE token was released around Meta’s rebranding in November 2021, after the latter provoked a sudden price increase in Decentraland’s MANA (fourfold in just a couple of days). This was a considerable surge, even taking into account the already upward trend of MATIC in the wider Polygon ecosystem. However, the collapse of several interconnected cryptocurrencies the following May marked the onset of the 2022 cryptowinter OECD (2022), which affected a large share of blockchain projects and lasted until mid 2023.
@@
Figure 4: Cryptocurrency price in USD during the two-year period of study for: ICE by Decentral Games, MANA by Decentraland, and MATIC by Polygon.
@@
For data on DG wearable upgrades, we again used Alchemy to retrieve the log events for the period of study of DG’s smart contract on Polygon [^2] Address 0xC9a67eD1472A76d064C826B54c144Ca00DAE4015 which manages the respective on-chain transactions, called “iceRegistrant”. We then used the aforementioned wearable metadata to derive the ranking level and rarity limit of the upgraded tokens.
@@
For in-world user tracking, we utilized the corresponding data archives from Atlas Data Corp, a blockchain analytics company that obtained a grant  [^3] https://decentraland.org/governance/project/?id=7422a99e-8a25-4625-9b30-ab688de5dade from Decentraland’s DAO to finance a system to keep record of user sessions. These data are log events (created every circa 20 seconds) of the current location of every connected user on the several servers that manage the 3D rendering of in-world scenes and avatars on Decentraland. Given that this system started operating in mid July 2022, we narrowed our analysis of in-world visits to only 10 months, from August 2022 to May 2023. Moreover, we take into consideration only users with an Ethereum-compatible address, that is, we discard the small minority of guest users not associated with an account. We then transformed the more than 677 million log events for this time frame into sessions with a start and end for each land parcel and account address. Finally, we segmented the respective sessions’ duration on a daily basis, adding 10 seconds to each to compensate for possible shortenings due to the 20-second polling rate of the log events.
5 Analyses and Results
5.1 RQ1: Overall wearable transfers
@@
From June 2021 to May 2023, a total of 4,629 collections with 8,450 wearable items (i.e., designs) were published by 1,517 distinct creators. Of these, we identified and manually verified two accounts [^4] Addresses 0x17a253c2ac0d5ba92cadbbf665e3390c9913dc5d and 0x00e5d44f6a296c10f159486f842838bd68f13e32 that created all of the DG wearables, based on marketplace offerings, item names, and user profiles on Decentraland. Both DG accounts (0.1% of creators) published 117 collections with 723 items (8.6% of the total available items). Henceforth, we refer to these as DG wearables.
@@
In our analysis on a monthly basis, we noticed great variation in the share of DG wearables over time, but interestingly, the variations are different for transfers and sales. On one hand, and as illustrated in Figure 5, the total volume of transfers reached its highest values in March and April of 2022, which coincides with the first edition of the Metaverse Fashion Week at the end of March. In those two months the share of DG wearables was only 3.3% and 4.5%, respectively. However, the transfers monthly share of DG wearables increased noticeably from September 2022 onward, reaching its peak in May 2023 with 85.3%, while the rest of wearable transfers decreased in number. This DG share increase is mostly related to a few wearable items with a common rarity limit and which give access only to ICE Poker Arcade, whose app was released precisely at the end of September. On the other hand, for the monthly sales volume in USD (illustrated in Figure 6), the highest monthly value for both total volume and share of DG wearables (62.9%) is November 2021, the month in which the first DG collections focused on ICE Poker were released for minting. Moreover, the timing also coincides with the wake of Meta’s rebranding, which provoked a surge and sustained high price in both ICE and MANA for that month, as shown in Figure 4.
Figure 5: Monthly number of transfers of Decentraland wearables on the Polygon blockchain and respective share associated with Decentral Games (DG). The upward trend in the bottom panel is related to wearables that grant access to the ICE Poker Arcade app, released in September 2022.
Figure 6: Monthly sales volume in USD of Decentraland wearables —made via the marketplaces of Decentraland and OpenSea— and the respective share associated with Decentral Games (DG). The bottom panel’s spike in November 2021 is related to the release of ICE-themed DG wearables and the high valuation of ICE in the wake of Meta’s rebranding.
@@
5.2 RQ2: In-world visits of poker casinos

### v3

@@ A.   Fashion in Virtual Worlds
Fashion in social virtual worlds has been a subject of research for years, especially in Second Life, which has been active since 2003 and introduced the concept of customizable virtual spaces and avatars [3]. Despite its recent decline, Second Life maintained for years a robust economy centered on avatar customization, with many users becoming in-world fashion designers [6, 8].
@@
Recently, the fashion industry’s attention has shifted towards popular MMO games, particularly through the creation of branded skins, with an increasing interest in newer blockchain-based virtual worlds [46]. Notably, several prestigious luxury brands are venturing into NFT collectibles [32], participating in events such as the Metaverse Fashion Week or designing wearables for Decentraland [24], as well as retailing and engaging in other kinds of customer engagement within its social virtual world [23]. In fact, Decentraland’s wearables are mainly distributed for free to boost engagement in other cryptoasset or Metaverse initiatives such as virtual casinos, with only a minor portion of them (3.4%) being sold on the platform’s marketplace [51].
B.   Poker and Online Gambling
@@
The chance-based nature of the game of poker appeals to a wide audience of casual players, while the element of skill, particularly evident over many hands, entices professional players [18]. The introduction of the Internet to gambling activities has provided easy access for the general public, round-the-clock access, shorter intervals between bets, and instant reinforcements [25]. After 2003, online poker experienced tremendous growth, with a total market value reaching 3 to 4 billion USD by 2010, with approximately 5.5 million players participating in virtual poker rooms across the globe [17]. Among such players, a subset has emerged as professionals, especially adept at multitabling —simultaneously playing multiple games to further reduce the element of chance. As reliance on poker earnings grew, these players developed a strong extrinsic motivation to play, even if not entirely removed from the gameplay. Further, it has been found that playing for real money strongly influences both game dynamics and player experiences in poker games [55].
@@
Notably, professional gamblers share psychological traits with problem gamblers, highlighting the complex interplay between skill and risk [40]. Problem gamblers, often drawn to low-stakes gambling, may develop distorted risk perceptions and ultimately underestimate financial risks in general [2]. Interestingly, players are socially motivated to gamble by the virtual communities they form, while these same communities also serve as safeguards against excessive behavior [49, 47]. Online poker’s unique subculture —which includes websites, forums, and chat rooms from all around the world— fosters cooperative behavior among players through the sharing of information and strategies. Nonetheless, the game’s inherent competitiveness persists and is reflected in the hierarchical structures within these global communities [44].
C.   Crypto Asset Trading
@@
Cryptocurrency has emerged as a disruptive force in the financial landscape since Bitcoin’s introduction in 2008. Nonetheless, its trading remains characterized by high volatility, often leading to substantial crashes followed by significant rebounds [16, 42]. NFTs entered the scene in 2014, gaining widespread attention in early 2021 due to very high-profile sales and causing considerable hype [30]. NFT buyers exhibit a profound enthusiasm for the foundational technologies that facilitate unique avenues for creative expression and the development of innovative content creation business models. However, these positive sentiments are counterbalanced by the contentiousness of digital ownership, prevalence of low-quality NFTs, risk of scams, potential money laundering activities, and evolving legal framework [48].
@@
There is also an overlap of personality and demographic traits among problem gamblers and cryptocurrency traders [31]. In both cryptocurrency trading and gambling, decisions are often based on limited information, short-term motives for gain, and with high uncertainty; hence, people attracted to the latter are likely to be attracted to the former [11]. Numerous operators in the crypto gambling sector provide access to a variety of betting options, games, and online casinos. However, these platforms often pose a significant risk as they are easily accessible to minors and vulnerable individuals, and typically offer minimal to no consumer protection. Further, they still remain largely unregulated in many countries [1].
D.   Play-to-Earn Games
@@
Blockchain technology has also boosted a form of online gaming known as “Play-to-Earn” (P2E), where players receive monetary rewards for their gameplay, mainly in the form of the game’s own cryptocurrency. Axie Infinity —a Pokémon-style battle game with collectible axolotl-like creatures and in-game tokenomics— is arguably one of the most well-known examples. Its popularity peaked in 2021, particularly in the Philippines and later Venezuela, where it represented a viable source of earnings for low-income people during the height of the Covid-19 pandemic [10]. However, it is known that extrinsically motivated gameplay, i.e., mainly driven by external rewards such as financial gain, may reduce intrinsic motivation and consequently have negative implications for player experience [12, 29]. In fact, a survey conducted with a group of Axie Infinity players in 2020–21 has shown that, while some players enjoyed social interactions, overall intrinsic motivation tended to be low, causing high amounts of stress and unwanted tension to reach daily quotas [10]. This assessment was later corroborated by a study that analyzed online conversations about the game [12].
@@
Interestingly, part of the initial success of Axie Infinity was due to delegation models that emerged organically within the community, called scholarships [10]. These allow players to share their knowledge or delegate their gameplay (and assets) to others for a share of the rewards. Nonetheless, while P2E games offer players from less-advantaged countries earning prospects above the national average wage, these might find themselves in exploitative schemes to provide low-cost labor for wealthier owners of NFTs [12]. In lieu of the community-based scholarships of Axie Infinity, ICE Poker integrates elements of P2E games by design, allowing players to earn tokens just by participating in certain events, and providing them with sanctioned asset delegation. Wealthy owners are thus offered an official mechanism to delegate their DG wearables to willing players for whom the initial investment is perhaps prohibitively expensive.
IV.   Methods
For our study, we focus on the two years after the launch of Decentraland wearables on Polygon, that is, from June 2021 to May 2023. For this, we undertake a data-driven approach based on public data available from blockchains and related datasets, focusing on obtaining macro-level patterns of actual user behavior related to our RQs. In the following, we describe the overall approach to answer our RQs and how we performed data collection.
A.   Overall approach
@@
For RQ1 (wearable transfers) we first retrieved all wearables created during the period of study, marking those created by DG based on metadata from Decentraland. We then retrieved from the Polygon blockchain the respective transfers, as well as sales and wearable upgrades from related sources. We then analyzed the proportion between DG wearables and the rest in terms of the number of transfers and sales volume (both in number and USD).
@@
For RQ2 (in-world visits to ICE poker casinos), we first collected in-world tracking data of Decentraland users, available only from 12 July 2022. Hence, we analyzed the last 10 months of the two-year period of interest, in terms of the number of accounts and time spent visiting the parcels in which ICE poker casinos are located, compared to the rest of Decentraland’s map.
@@
Our descriptive and statistical analyses mostly concern proportions, i.e., the share of wearables or land associated with DG compared with the rest of Decentraland. To test for significant differences in proportions we used two-sided tests with significance level $\alpha=0.05$. When analyzing samples we used $\chi^{2}$ tests, either goodness-of-fit test or association test according to the case; when analyzing populations we used a binomial test; for both kinds of test we use relative risk (RR) to measure effect sizes.
B.   Data collection
@@
For wearable metadata (e.g., collection, category, creator), we used several of the openly available Decentraland’s query endpoints on TheGraph, a service for querying blockchain networks, and whose definitions are open source. [^1] https://github.com/decentraland For transfers, we collected the respective transactions from Polygon via the blockchain web service Alchemy. For sales, we collected data from two different sources: 1) Decentraland’s Marketplace for internal sales, using the aforementioned TheGraph query endpoints; and 2) OpenSea, the largest overall NFT marketplace and main means for external sales, as well as the initial default marketplace for DG wearables.
@@
We also retrieved from Yahoo! Finance the price of the few cryptocurrencies used in the collected wearable sales, which we adjusted for inflation. Price changes of these cryptocurrencies also serve to gauge interest on the studied platforms. As can be seen in Figure IV, the ICE token was released around the rebranding of Facebook into Meta in November 2021, with the latter provoking a sudden price increase in Decentraland’s MANA (fourfold in just a couple of days). This was a considerable surge, even taking into account the already upward trend of MATIC, Polygon’s cryptocurrency. However, the collapse of several interconnected cryptocurrencies the following May marked the onset of the 2022 cryptowinter [42], which affected a large share of blockchain projects and lasted until mid 2023.
@@
Figure IV: Cryptocurrency price in USD during the two-year period of study for: ICE by Decentral Games, MANA by Decentraland, and MATIC by Polygon.
@@
For data on wearable upgrades, we again used Alchemy to retrieve the log events of the DG’s smart contract on Polygon [^2] Address 0xC9a67eD1472A76d064C826B54c144Ca00DAE4015 that managed such transactions. We then used the aforementioned wearable metadata to derive the ranking level and rarity limit of the upgraded tokens.
@@
For in-world user tracking, we utilized the corresponding data archives from Atlas Data Corp, a blockchain analytics company that obtained a grant  [^3] https://decentraland.org/governance/project/?id=7422a99e-8a25-4625-9b30-ab688de5dade from Decentraland’s DAO to keep record of user sessions. These data are log events (created every circa 20 seconds) of the current location of every connected user on the several servers that manage the 3D rendering of in-world scenes and avatars on Decentraland. Given that this system started operating in mid July 2022, we narrowed our analysis of in-world visits to only 10 months, from August 2022 to May 2023. Moreover, we take into consideration only users with an Ethereum-compatible address, that is, we discard the small minority of guests not associated with an account. We then transformed the more than 677 million log events for this time frame into sessions with a start and end for each land parcel and account address. Finally, we segmented the respective sessions’ duration on a daily basis, adding 10 seconds to each to compensate for possible shortenings due to the 20-second polling rate of the log events.
V.   Analyses and Results
A.   RQ1: Overall wearable transfers
@@
From June 2021 to May 2023, a total of 4,629 collections with 8,450 wearable items (i.e., designs) were published by 1,517 distinct creators. Of these, we identified and manually verified two accounts [^4] Addresses 0x17a253c2ac0d5ba92cadbbf665e3390c9913dc5d and 0x00e5d44f6a296c10f159486f842838bd68f13e32 that created all of DG wearables, based on marketplace offerings, item names, and user profiles on Decentraland. Both DG accounts (0.1% of creators) published 117 collections with 723 items (8.6% of the total available items). Henceforth, we refer to these as DG wearables.
@@
In our analysis on a monthly basis, we noticed great variation in the share of DG wearables over time, but in different ways for transfers and sales. On one hand, and as illustrated in Figure V, the total volume of transfers reached its highest values in March and April of 2022, which coincides with the first edition of the Metaverse Fashion Week at the end of March. In those two months the share of DG wearables was only 3.3% and 4.5%, respectively. However, the transfers monthly share of DG wearables increased noticeably from September 2022 onward, reaching its peak in May 2023 with 85.3%, while the rest of wearable transfers decreased in number. This DG share increase is mostly related to a few wearable items with a common rarity limit and which give access only to ICE Poker Arcade, whose app was released precisely at the end of September. On the other hand, for the monthly sales volume in USD (illustrated in Figure VI), the highest monthly value for both total volume and share of DG wearables (62.9%) is November 2021, the month in which the first DG collections focused on ICE Poker were released for minting. Moreover, the timing also coincides with the wake of Meta’s rebranding, which provoked a surge and sustained high price in both ICE and MANA for that month, as shown in Figure IV.
@@
Figure V: Monthly number of transfers of Decentraland wearables on the Polygon blockchain and the respective Decentral Games (DG) share.
Figure VI: Monthly sales volume in USD of Decentraland wearables —made via the marketplaces of Decentraland and OpenSea— and the respective share associated with Decentral Games (DG).
B.   RQ2: In-world visits of poker casinos